kim had $1.25 in nickels and dimes in her purse. if the the number of her nickels is four more than the number of dimes, how many nickels does she have?
Answer by checkley77(12844)

N=D+4
.10D+.05N=1.25
.10D+.05(D+4)=1.25
.10D+.05D+.20=1.25
.15D=1.25-20
.15D=1.05
D=1.05/.15
D=7 NUMBER OF DIMES.
N=7+4=1 NUMBER OF NICKELS.
PROOF:
.10*7+.05*11=1.25
.70+.55=1.25
1.25=1.25